# Equal Contribution   Equal Advising https://sirui-xu.github.io/InterDreamer/

### v1

@@ 3.2 Low-Level Control
With the information $g$ derived from the description $p$, the low-level control aims to create an initial state $\boldsymbol{s}_{1}$ and a sequence of human actions $\{\boldsymbol{a}_{t}\}_{t=1}^{T}$, such that they correspond with the objectives outlined by $g$. In this section, we detail the text-to-motion model and interaction retrieval system designed to fulfill the task.
Text-to-Motion. We construct the text-to-motion model $\pi$ designed to develop actions to be executed in the world model. At each timestep $t$, $\pi$ receives the sequence of previous actions $\{\boldsymbol{a}_{i}\}_{i=1}^{t}$ and the text tokens extracted from $g=L(p)$, and produces a preliminary next action ${\boldsymbol{a}_{t+1}}$, which will be adjusted to the ultimate action $\boldsymbol{a}_{t+1}^{\ast}$ through an optimization process that intertwines it with the object state, introduced in Sec. 3.4. Thus, the overall process coupled with the optimization can be formally defined as $\boldsymbol{a}_{t+1}\sim\pi(\boldsymbol{a}_{t+1}|\boldsymbol{s}_{t},\{\boldsymbol{a}_{i}\}_{i=1}^{t},g)$. The initial action $\boldsymbol{a}_{1}\sim\pi(\boldsymbol{a}_{1}|g)$ is influenced by the text from $g$ without prior actions or states, which will be used in the interaction retrieval. As detailed in Sec. 4.1, $\pi$ is capable of leveraging existing text-to-motion models, including MDM [93], MotionDiffuse [130], ReMoDiffuse [131], and MotionGPT [35].
Interaction Retrieval. The interaction retrieval component $R$ sets up an initial state $\boldsymbol{s}_{1}\sim R(\boldsymbol{a}_{1},g)$, given the initial action $\boldsymbol{a}_{1}$ generated by the text-to-motion model. We present here a user-friendly pipeline that is based on handcraft rules and a learning-based approach in Sec. B.1 of supplementary. First, we build a database by collecting HOI frames from the target dataset, e.g., the BEHAVE [7] or CHAIRS [36] dataset. The indexing key for retrieval is a tuple consisting of the body part in contact and the category of the object involved. The value for retrieval is a per-frame contact map, i.e., a list of $K$ vertex pairs $\{(d_{h}^{i},d_{o}^{i})\}_{i=1}^{K}$. Here, $d_{h}^{i}$ indexes the contact vertex on the human’s surface, while $d_{o}^{i}$ indexes the corresponding contact vertex on the object’s surface. This contact map is linked with their key, thus establishing a searchable record of interactions. During the inference stage, equipped with the body part and object information provided by the high-level planning (Sec. 3.1), we utilize them as a key to retrieve all relevant contact maps from our database. We sample one map $\{(d_{h}^{i},d_{o}^{i})\}_{i=1}^{K}$ according to a pre-defined metric, and employ it to ascertain an object state $\boldsymbol{s}_{1}\sim R(\boldsymbol{a}_{1},g)$, harmonious with the initial human action $\boldsymbol{a}_{1}$, as the initiation of the interaction. More details are provided in Sec. B.1 of supplementary.
@@ 3.3 World Model
Our world model combines a dynamics model and the optimization process, dedicated to simulating state transitions affected by applied actions. While drawing inspiration from similar concepts utilized in robotics [108, 83] and autonomous driving systems [43], we use it here for rollout without model-based learning. This model, trained on the 3D HOI dataset, serves a role similar to a physics simulator but is much simpler – it takes the preceding object state $\boldsymbol{s}_{t}$ along with a pair of consecutive actions $\boldsymbol{a}_{t}$ and $\boldsymbol{a}_{t+1}$, and predicts the subsequent object state $\boldsymbol{s}_{t+1}$. The interplay between the low-level control and the world model ultimately produces a coherent interaction rollout.
@@
In designing the dynamics model, a naïve method would be directly taking the raw actions and past state as its input. This method, however, suffers from a severe generalization problem during inference: the dynamics model is likely to encounter some human actions that do not exist in the training set, since our text-to-motion model is not trained on the HOI dataset. We provide ablation studies as evidence for this claim in Sec. 4. Thus, instead of directly modeling the complex distribution of interactions, we approach the interactions from the contact vertices on the object, as shown in Fig. 2. This locality ensures that the dynamics model remains focused on interactions in the contact region, without being distracted by the motion of body parts that are irrelevant to the object.
Input Representation. Specifically, at each time step $t$, we abstract the past action as $H$ historical vertex trajectories $\{\{\boldsymbol{v}_{i}^{j}\}_{j=1}^{N}\}_{i=1}^{H}$, and the future action as $F$ future vertex trajectories $\{\{\boldsymbol{v}_{i}^{j}\}_{j=1}^{N}\}_{i=H+1}^{H+F}$, where non-fixed variable $N$ is the number of sampled contact vertices. Note that we train our dynamics model to forecast over a longer duration than the past motion ($F>H$), while only the foremost future action will be used for autoregressive generation during the inference, to facilitate the long-term generation process, as suggested in [16].
To determine these $N$ vertices, we start with object’s signed distance fields $\{\mathbf{sdf}_{i}\}_{i=1}^{H}$ over the past $H$ frames, derived from the past state $\boldsymbol{s}_{t}$. We then sample vertices that meet the following criteria:
$|\mathbf{sdf}_{i}(\boldsymbol{v}_{i}^{j})|\leq\delta_{1},\quad\forall i=1,\ldots,H,\quad\forall j,$ (1)
$\|\boldsymbol{v}_{i}^{j}-\boldsymbol{v}_{i}^{k}\|\geq\delta_{2},\quad\forall j\neq k,$ (2)
where $\delta_{1}$ and $\delta_{2}$ are two hyperparameters. The objective is to sparsely sample contact vertices, while ensuring they are sufficient to encompass the interaction.
We characterize each vertex trajectory $\{\boldsymbol{v}_{i}^{j}\}_{i=1}^{H+F}$ with a feature $\boldsymbol{f}^{j}$ to provide information in addition to motion, which includes (i) vertex coordinates at T-pose, providing information about the position of the human vertex on the body surface; (ii) the vertex-to-object surface distance, indicating vertex’s impact on the object; and (iii) the vertex’s velocity relative to its nearest object vertex.
Architecture. An overview of the architecture is demonstrated in Fig. 2. An unconditional dynamics block can be initiated as $\mathcal{G}(\boldsymbol{x}_{k},\boldsymbol{\Theta})$, mapping the input feature map $\boldsymbol{x}_{k}$ at the $k$-th layer to another feature map, with $\Theta$ denoting the dynamics model parameters. To incorporate human vertex controls, we introduce a secondary network $\mathcal{F}(\boldsymbol{y}^{j}_{k},\boldsymbol{\Theta}_{v})$ operating on $N$ vertex features $\{\boldsymbol{y}^{j}_{k}\}_{j=1}^{N}$, where $\boldsymbol{\Theta}_{v}$ is its parameters. With a cross-attention layer $\mathrm{Attn}$, the final dynamics block is formulated as:
$\boldsymbol{x}_{k+1},\{\boldsymbol{y}^{j}_{k+1}\}_{j=1}^{N}=\mathrm{Attn}(\mathcal{G}(\boldsymbol{x}_{k},\boldsymbol{\Theta}),\{\mathcal{F}(\boldsymbol{y}^{j}_{k},\boldsymbol{\Theta}_{v})\}_{j=1}^{N}).$ (3)
@@
We stack multiple blocks to form the dynamics model. The initial input, $\boldsymbol{x}_{0}$, corresponds to the previous state $\boldsymbol{s}_{t}$, while each $\boldsymbol{y}^{j}_{0}$ represents the feature of the vertex trajectory, containing both the trajectory $\{\boldsymbol{v}_{i}^{j}\}_{i=1}^{H+F}$ and its associated feature vector $\boldsymbol{f}^{j}$. The output of this model, $\boldsymbol{s}_{t+1}$, is preliminary and subject to further optimization as detailed in Sec. 3.4, which will yield the final future state $\boldsymbol{s}_{t+1}^{\ast}$. We utilize the Mean Squared Error loss to train the dynamics model. For more details, please refer to Sec. B.2 of supplementary.
@@ 3.4 Optimization
Optimization serves as a role to introduce prior knowledge and avoid the accumulation of errors. During inference, we input the initial action $\boldsymbol{a}_{t+1}$ and state $\boldsymbol{s}_{t+1}$ and refine them into the fine-grained action $\boldsymbol{a}_{t+1}^{\ast}$ and state $\boldsymbol{s}_{t+1}^{\ast}$. This refinement is achieved through gradient descent on the human and object pose parameters. Our optimization includes several loss terms: a fitting loss to align $\boldsymbol{a}_{t+1}^{\ast}$ and $\boldsymbol{s}_{t+1}^{\ast}$ with their preliminary counterparts, a velocity loss for temporal smoothness, a contact loss to promote occurring contacts, and a collision loss to reduce penetration in the interaction. For efficiency, we perform optimization only if the loss is above a threshold. Specifically, given the reference interaction sequence $\{\boldsymbol{h}_{i}\}_{i=1}^{L}$ and $\{\boldsymbol{o}_{i}\}_{i=1}^{L}$ of arbitrary length $L$, derived from previous steps, we apply gradient descent to optimize human pose sequence $\{\boldsymbol{h}_{i}^{\ast}\}_{i=1}^{L}$ and object pose sequence $\{\boldsymbol{o}_{i}^{\ast}\}_{i=1}^{L}$, using the loss function,
$E_{\mathrm{fit}}=\sum_{i=1}^{L}(\|\boldsymbol{h}_{i}^{\ast}-\boldsymbol{h}_{i}\|_{1}+\|\boldsymbol{o}_{i}^{\ast}-\boldsymbol{o}_{i}\|_{1}).$ (5)
$E_{\mathrm{vel}}=\sum_{i=1}^{L-1}(\|\boldsymbol{h}_{i+1}^{\ast}-\boldsymbol{h}_{i}^{\ast}\|_{1}+\|\boldsymbol{o}_{i+1}^{\ast}-\boldsymbol{o}_{i}^{\ast}\|_{1}).$ (6)
Contact Loss. We leverage a contact loss to encourage body part to contact the object surface, if they are close to each other in the initial interaction,
$E_{\mathrm{cont}}=\sum_{i=1}^{L}\sum_{d_{h}\in\mathcal{T}_{i}}\min_{d_{o}}\|\boldsymbol{v}_{\boldsymbol{o}_{i}^{\ast}}[d_{o}]-\boldsymbol{v}_{\boldsymbol{h}_{i}^{\ast}}[d_{h}]\|_{2}.$ (7)
where $\boldsymbol{v}_{\boldsymbol{h}_{i}^{\ast}}[d_{h}]$ denotes the vertex on the human body surface, and $\boldsymbol{v}_{\boldsymbol{o}_{i}^{\ast}}[d_{o}]$ represents the corresponding vertex on the surface of the object. And $\mathcal{T}_{i}=\{d_{h}|\min_{d_{o}}\\ \|\boldsymbol{v}_{\boldsymbol{o}_{i}}[d_{o}]-\boldsymbol{v}_{\boldsymbol{h}_{i}}[d_{h}]\|_{2}\leq\epsilon\}$ includes the index of reference human vertex $\boldsymbol{v}_{\boldsymbol{h}_{i}}[d_{h}]$ that is close to the reference object vertex $\boldsymbol{v}_{\boldsymbol{o}_{i}}[d_{o}]$, where $\epsilon$ is a hyperparameter, $d_{h}$ and $d_{o}$ are vertex indices for human mesh and object mesh, respectively.
Penetration Loss. Given the signed-distance field of the human pose $\textbf{sdf}_{\boldsymbol{h}_{i}^{\ast}}$, we employ a penetration loss to penalize the body-object interpenetration,
$E_{\mathrm{pene}}=-\sum_{i=1}^{L}\sum_{d_{o}}\min(\textbf{sdf}_{\boldsymbol{h}_{i}^{\ast}}(\boldsymbol{v}_{\boldsymbol{o}_{i}^{\ast}}[d_{o}]),0).$ (8)
Figure 3: Qualitative results on the BEHAVE dataset [7]. The interaction sequences are presented through a time-series visualization where color changes denote progression through frames. Frames are separately visualized when the pelvis remains nearly static. Here, our synergized knowledge comes from GPT-4 [67] and MotionGPT [35].
4 Experiments
Extensive comparisons evaluate the performance of our InterDreamer across two motion-relevant tasks. Details of the evaluation settings are provided in Sec. 4.1. We present both quantitative (Sec. 4.2) and qualitative results (Sec. 4.3) for our approach. Additionally, we perform ablation studies to verify the efficacy of each component within our framework. These studies also cover the interaction prediction task [116] to evaluate our dynamics model. Additional details and results are presented Sec. C and Sec. D of the supplementary.
Figure 4: Qualitative results in more challenge scenarios with free-form input not from our annotations, showing the ability of our InterDreamer to fit different object sizes and handle complex and long sequences. Here, our synergized knowledge comes from GPT-4 [67] and MotionGPT [35].
Table 1: Quantitative results on human motion quality on the BEHAVE dataset with our annotation. We show that our high-level planning narrows the distributional gap and effectively adapts single human generators into zero-shot human-object interaction generation. To evaluate R-Precision, a batch size of 16 is selected.
Methods Planning (Ours) R-Precision FID MM Dist Multimodality Diversity
@@ 4.1 Experimental Setup
Datasets. We evaluate our model on the BEHAVE dataset [7], which includes recordings of 8 individuals interacting with 20 everyday objects. Our analysis focuses on 18 objects for which interaction sequences are available at 30 Hz. The human pose is modeled using SMPL-H [58, 82], with hand poses set to an average pose due to the absence of detailed hand pose in the dataset. Object poses are rotation matrices and translations. We manually segment the long interaction sequences in the test set, and annotate them with descriptions as well as their starting and ending indices, leading to $532$ subsequences for evaluation. The CHAIRS [36] dataset encompasses the capture of 46 subjects, represented via SMPL-X [70] bodies, engaging with 81 distinct types of chairs and sofas. Without whole dataset annotations, we use the CHAIRS dataset solely for qualitative evaluation. We further use GPT-4 [67] to rephrase and diversify annotations: 1) less complexity: someone holds a backpack and steps left; 2) our annotation: a person holds a backpack in front of them with both hands and takes a step to the left; 3) more complexity: with both hands, a person clutches a heavy backpack firmly and brings it close to their body, then steps to the left with their left leg.
@@
Figure 5: Qualitative results on the CHAIRS dataset [36]. Our dynamics model trained on the BEHAVE dataset [7] generalizes well on the CHAIRS dataset unseen in training. Interaction sequences are visualized through a time-series style where color changes denote progression through frames. Frames are separately visualized. Here, high-level planning and low-level control use GPT-4 [67] and MotionGPT [35], respectively.
Table 2: Quantitative results on the zero-shot text-to-interaction task (left) and on the interaction prediction task (right). Our dynamics model with vertex-based control generates interactions with the best quality.
Methods Zero-shot text-to-interaction Interaction prediction [116]
w/o control 0.424 533 123 256 228
w/ marker control (InterDiff [116]) 0.219 484 123 226 164
w/ raw control 0.325 957 129 265 218
w/ vertex control (ours) 0.151 443 119 221 156
Metrics. The evaluation metrics are divided into three categories: (i) Human motion quality: The Fréchet Inception Distance (FID) measures the distributional distance between the generated motions and ground truth. The MultiModality (Multimodality) and Diversity metrics assess the variance in generated human motion. R-Precision evaluates the consistency between the text and the generated human motion within the latent feature space. MultiModal distance (MM Dist) is the distance between the motion feature and the text feature. We follow [25] to generate motion and text features. (ii) Interaction quality: We propose a metric to measure the distance between contact maps of real interactions and those generated (CMD). The per-sequence contact map is defined by the percentage of time that each body part is actively in contact. The detailed formulation is provided in Sec. C of supplementary. We measure the collision (Pene. [116]), which calculates the average percentage of object vertices that have non-negative values in the human signed distance fields [69]. (iii) Object motion accuracy: The dynamics model’s performance in the interaction prediction task [116] is evaluated by the accuracy of predicted object motions, including Trans. Err., the average distance between predicted and ground truth, and Rot. Err., the average distance between the predicted and ground truth.
@@
Figure 6: (a) Ablation study on the high-level planning. On the left are results from MotionGPT [35] based on raw descriptions (“w/o planning”); on the right are results with our planning (“w/ planning”). Free-form descriptions with the out-of-distribution object name lead to results that bear little resemblance to the description. (b) We visualize CLIP [78] features of descriptions on HumanML3D [25], our initial raw annotations (“w/o planning”), and the annotations processed through our high-level planning framework (“w/ planning”). The CLIP features are processed via t-SNE [61].

### v2

@@ 3.2 Low-Level Control
With the information $g$ derived from the description $p$, the low-level control aims to create a sequence of human actions $\{\bm{a}_{t}\}_{t=1}^{T}$ by a text-to-motion model, and an initial state $\bm{s}_{1}$ by interaction retrieval, such that they correspond to the objectives outlined by $g$.
Text-to-Motion. We utilize a text-to-motion model $\pi$ to develop actions to be executed in the world model. At each timestep $t$, $\pi$ receives the sequence of previous actions $\{\bm{a}_{i}\}_{i=1}^{t}$ and the text tokens encoded from the rewritten description in $g=L(p)$, and produces a next action $\bm{a}_{t+1}$}, which later in Sec. 3.4 will be adjusted through an optimization process that intertwines actions with the object state $\bm{s}_{t}$. Thus, the overall process can be formally defined as $\bm{a}_{t+1}\sim\pi(\bm{a}_{t+1}|\bm{s}_{t},\{\bm{a}_{i}\}_{i=1}^{t},g)$, while the initial action $\bm{a}_{1}\sim\pi(\bm{a}_{1}|g)$ is influenced merely by context $g$ without prior actions or states, which will be used in interaction retrieval. $\pi$ builds upon existing text-to-motion models, where we evaluate MDM [117], MotionDiffuse [172], ReMoDiffuse [173], and MotionGPT [46].
Interaction Retrieval. The interaction retrieval component $R$ establishes the initial state $\bm{s}_{1}\sim R(\bm{s}_{1}|\bm{a}_{1},g)$, based on the initial action $\bm{a}_{1}$ generated by the text-to-motion model. We propose a user-friendly pipeline for this purpose built on handcrafted rules. First, we create databases by extracting HOI frames from the training sets of each target datasets — BEHAVE [7], OMOMO [66], and CHAIRS [47]. The indexing key for retrieval is a tuple consisting of the body part in contact and the category of the involved object. Each retrieval value is a per-frame contact map, represented by a list of $K$ vertex pairs $\{(d_{h}^{i},d_{o}^{i})\}_{i=1}^{K}$. Here, $d_{h}^{i}$ refers to the contact vertex on the human surface, while $d_{o}^{i}$ refers to the corresponding contact vertex on the object surface. This contact map is linked to its corresponding key, creating a searchable record of interactions. During the inference stage, using the body part and object information provided by the high-level planning (Sec. 3.1), we retrieve all relevant contact maps from the database. We then sample one map $\{(d_{h}^{i},d_{o}^{i})\}_{i=1}^{K}$ and use it to establish the object state $\bm{s}_{1}\sim R(\bm{s}_{1}|\bm{a}_{1},g)$, thus initializing the interaction. Further details including how we ensure consistency between the sampled state and human action are provided in Sec. B.1 of the Appendix. We also discuss an alternative learning-based approach in Sec. B.1.
@@ 3.3 World Model
Our world model combines a dynamics model and the optimization process, dedicated to simulating state transitions affected by applied actions. While drawing inspiration from similar concepts utilized in robotics [135, 102] and autonomous driving systems [55], we use it here to generate HOI trajectories. This model, trained on the training set of a 3D HOI dataset such as BEHAVE [7], serves a similar role as a simulator but is much simpler – it takes the preceding object state $\bm{s}_{t}$ along with a pair of consecutive actions $\bm{a}_{t}$ and $\bm{a}_{t+1}$, and predicts the subsequent object state $\bm{s}_{t+1}$. The interplay between the low-level control and the world model ultimately produces a coherent interaction rollout.
@@
In designing the dynamics model, a naïve method would be directly taking raw actions, states, and object geometry as input. However, this suffers from a severe generalization problem during inference: the dynamics model is likely to encounter human actions and object geometry that do not exist in the training set, since our text-to-motion model is not trained with object interaction data. To overcome this limitation, we instead focus on encoding interactions through the contact vertices on the object, which capture both the action and object geometry, as shown in Figure 2. This locality ensures that the dynamics model remains focused on interactions in the contact region, without being distracted by the motion of body parts and geometry details that are irrelevant to the interaction.
Input Representation. Specifically, at each timestep $t$, we abstract the past actions as $H$ historical vertex trajectories $\{\{\bm{v}_{i}^{j}\}_{j=1}^{N}\}_{i=1}^{H}$, and the future actions as $F=m$ future vertex trajectories $\{\{\bm{v}_{i}^{j}\}_{j=1}^{N}\}_{i=H+1}^{H+F}$, where non-fixed variable $N$ is the number of sampled contact vertices, and $m$ is the length of segments as mentioned in the overview of Sec 3. Note that we train our dynamics model to forecast over a longer duration than the past motion ($F>H$), only the foremost future action will be used for autoregressive generation during the inference, as suggested in [19]. To determine these $N$ vertices, we start with object’s signed distance fields $\{\mathbf{sdf}_{i}\}_{i=1}^{H}$ over the past $H$ frames, derived from the past state $\bm{s}_{t}$. We then sample vertices that meet the following criteria: $|\mathbf{sdf}_{i}(\bm{v}_{i}^{j})|\leq\delta_{1},\mathbf{sdf}_{i}(\bm{v}_{i}^{k})|\leq\delta_{1},\forall i=1,\ldots,H,\forall j$, and $\|\bm{v}_{i}^{j}-\bm{v}_{i}^{k}\|\geq\delta_{2},\forall j\neq k$, where $\delta_{1}$ and $\delta_{2}$ are two hyperparameters. The objective is to sparsely sample contact vertices while ensuring that they are sufficient to encompass the interaction. We characterize each vertex trajectory $\{\bm{v}_{i}^{j}\}_{i=1}^{H+F}$ with a feature $\bm{f}^{j}$ to provide (i) human vertex coordinates at T-pose, providing information about the position of the human vertex on the body surface; (ii) the vertex-to-object surface vector, indicating vertex’s impact on the object as well as inherently including information related to the object’s shape; and (iii) the vertex’s velocity relative to its nearest object vertex. Thus, the model needs to learn how the features of human action $\bm{f}^{j}$ affect the evolution of the state of the object.
@@
Architecture. As demonstrated in Figure 2, the network comprises two components: $\mathcal{G}$ that operates without contact vertex conditions, applicable in scenarios where no contact occurs, and $\mathcal{F}$, which incorporates contact vertex conditions into the object trajectory when contact is present. The k-th layer of $\mathcal{G}$ can be initiated as $\mathcal{G}_{k}(\bm{x}_{k},\bm{\Theta})$, mapping the input feature map $\bm{x}_{k}$ at the $k$-th layer to another feature map, with $\Theta$ denoting the MLP’s parameters. To incorporate human vertex controls, we introduce a second network $\mathcal{F}_{k}(\bm{y}^{j}_{k},\bm{\Theta}_{v})$ operating on $N$ vertex features $\{\bm{y}^{j}_{k}\}_{j=1}^{N}$, where $\bm{\Theta}_{v}$ is its parameters. With a cross-attention layer $\mathrm{Attn}$, a dynamics block is formulated as: $\bm{x}_{k+1},\{\bm{y}^{j}_{k+1}\}_{j=1}^{N}=\mathrm{Attn}(\mathcal{G}_{k}(\bm{x}_{k},\bm{\Theta}),\{\mathcal{F}_{k}(\bm{y}^{j}_{k},\bm{\Theta}_{v})\}_{j=1}^{N}).$ We stack multiple dynamics blocks to form the model. The initial input, $\bm{x}_{0}$, corresponds to the previous state $\bm{s}_{t}$, while each $\bm{y}^{j}_{0}$ represents the feature of the vertex trajectory, containing both the trajectory $\{\bm{v}_{i}^{j}\}_{i=1}^{H+F}$ and its associated feature vector $\bm{f}^{j}$. The output of this model is preliminary and subject to further optimization as introduced in Sec. 3.4, which will yield the final future state. We utilize the Mean Squared Error loss to train the dynamics model. For more explanations, please refer to Sec. B.2 of the Appendix.
@@ 3.4 Optimization
Optimization plays a role in introducing prior knowledge and avoiding the accumulation of errors. During inference, we input the action $\bm{a}_{t+1}$ and state $\bm{s}_{t+1}$ and refine them. This refinement is achieved through gradient descent on the human and object pose parameters. Our optimization includes several loss terms: a fitting loss to align optimized results with their preliminary one, a velocity loss for temporal smoothness, a contact loss to promote occurring contact, and a collision loss to reduce penetration. We provide detailed formulations in Sec. B.3 of the Appendix.
Table 1: Quantitative results on evaluating the dynamics model. Our dynamics model with vertex-based action generates interactions of the best quality.
Methods Text-to-Interaction Interaction Prediction [148]
w/o action 0.424 533 123 256 228
contact markers as action (InterDiff [148]) 0.219 484 123 226 164
human motion as action 0.325 957 129 265 218
contact vertices as action (ours) 0.151 443 119 221 156
Table 2: Quantitative results on human motion quality given our annotation on the BEHAVE [7] dataset. We show that our high-level planning effectively adapts single human generators into human-object interaction generation. To evaluate R-Precision, a batch size of 16 is selected.
Methods Planning (Ours) R-Precision FID $M$ Dist Multimodality Diversity
@@
Table 3: Quantitative results on human motion quality on the OMOMO [66] dataset with their provided annotation. We show that our high-level planning narrows the distribution gap and adapts single human generators into human-object interaction generation. To evaluate R-Precision, a batch size of 32 is selected.
Methods Planning (Ours) R-Precision FID $M$ Dist Multimodality Diversity
Top 1 Top 2 Top 3
Ground Truth - 0.044±0.004 0.095±0.008 0.151±0.009 0.000±0.000 6.858±0.006 - 5.660±0.110
MDM [117] $\times$ 0.056±0.005 0.096±0.007 0.135±0.006 16.638±0.631 7.110±0.063 2.446±0.456 5.862±0.520
$\checkmark$ 0.062±0.006 0.109±0.004 0.155±0.008 15.735±0.285 6.889±0.082 2.663±0.520 6.461±0.841
MotionDiffuse [172] $\times$ 0.048±0.006 0.094±0.008 0.143±0.013 15.442±0.231 5.799±0.054 1.658±0.209 5.981±0.516
$\checkmark$ 0.075±0.005 0.141±0.015 0.189±0.009 10.815±0.093 5.916±0.094 1.677±0.264 5.718±0.522
ReMoDiffuse [173] $\times$ 0.062±0.003 0.111±0.005 0.160±0.012 15.479±0.209 5.690±0.049 1.179±0.145 6.032±0.540
$\checkmark$ 0.067±0.004 0.127±0.006 0.174±0.006 14.560±0.080 5.678±0.033 1.193±0.202 5.368±0.417
MotionGPT [46] $\times$ 0.061±0.005 0.114±0.006 0.152±0.006 18.472±0.528 6.358±0.076 4.553±0.244 6.726±0.156
$\checkmark$ 0.064±0.007 0.121±0.007 0.164±0.009 17.512±0.498 6.287±0.041 4.470±0.191 7.048±0.169
Figure 3: Qualitative results on free-form text input. The interaction sequences, with textures from [13], are presented through a time-series visualization.
4 Experiments
Extensive comparisons evaluate the performance of our InterDreamer across two motion-relevant tasks. Details of the evaluation settings are provided in Sec. 4.1. We present both quantitative (Sec. 4.2) and qualitative (Sec. 4.3) results for our approach. Additionally, we perform ablation studies to verify the efficacy of each component within our framework. These studies also cover the interaction prediction task [148] to evaluate our dynamics model. Additional details and results are presented in Sec. C and Sec. D of the Appendix. Please refer to our website for video results.
Figure 4: Qualitative results in more challenge scenarios with free-form input not from our annotations, showing the ability of our InterDreamer to fit object sizes and handle complex and long sequences. Here, our synergized models are GPT-4 [87] and MotionGPT [46].
@@ 4.1 Experimental Setup
Datasets. We use BEHAVE [7], CHAIRS [47], and OMOMO [66] datasets for quantitative evaluation. The BEHAVE dataset includes recordings of 8 individuals interacting with 20 everyday objects, and our analysis focuses on 18 objects for which interaction sequences are available at 30 Hz. The human pose is modeled using SMPL-H [101], with hand poses set to an average pose due to the absence of detailed hand pose in the dataset. We manually segment the long interaction sequences in the test set, and annotate them with descriptions as well as their starting and ending indices, leading to $532$ subsequences for evaluation. For qualitative evaluation, we go beyond using annotations specifically created and employ free – form text to demonstrate our model’s capability on out-of-distribution inputs. To assess our model’s performance with novel objects, we expand our retrieval database to include objects from the OMOMO [66] and CHAIRS [47] datasets, while we do not fine-tune the dynamics model on them–a direct qualitative evaluation without additional adaptation.
@@
Figure 5: Qualitative results on the CHAIRS [47] dataset. Our dynamics model trained on the BEHAVE [7] dataset generalizes well on the CHAIRS objects unseen in training. Frames are separately visualized. Here, our synergized models are GPT-4 [87] and MotionGPT [46].
Figure 6: Results from the interaction retrieval. We demonstrate that our proposed retrieval approach based on handcraft rules can extract diverse and realistic interactions.
@@
Figure 7: (a) Ablation study on the high-level planning. On the left are results from MotionGPT [46] using free-form descriptions, and on the right are results with our planning additionally. Without planning, the motion generative model struggles to interpret free-form HOI descriptions and generate semantically-aligned motion. (b) We visualize CLIP [97] features of text on HumanML3D [34] via t-SNE [121], raw HOI descriptions (“w/o planning”), and HOI descriptions processed through our high-level planning (“w/ planning”). See Table 5 for quantitative measurements.
Metrics. The evaluation metrics are divided into three categories: (i) Human motion quality: The Fréchet Inception Distance (FID) measures the distance between the generated motion and ground truth. The MultiModality (Multimodality) and Diversity metrics assess the variance in generated human motion. R-Precision evaluates the consistency between the text and the generated human motion within the latent space. MultiModal distance (MM Dist) is the distance between the motion feature and the text feature. We follow [34] to generate motion and text features. (ii) Interaction quality: We propose CMD to measure the distance between contact maps of real interactions and those generated. The per-sequence contact map is defined by the percentage of time that each body part is actively in contact. The detailed formulation is provided in Sec. C of the Appendix. We also measure the collision (Pene. [148]), which calculates the average percentage of object vertices that have non-negative values in the human signed distance fields [89]. (iii) Object motion accuracy: The dynamics model’s performance in the interaction prediction task [148] is evaluated by the accuracy of predicted object motion, including Trans. Err., the average distance between predicted and ground truth, and Rot. Err., the average distance between the predicted and ground truth.
Baselines. Most recent research on text-to-HOI synthesis follows a supervised learning approach [90, 28], making direct quantitative comparisons unfair. Therefore, we primarily focus on qualitative comparisons with these methods. To enable quantitative evaluation, we develop a range of baselines to assess both the overall performance of our framework and the effectiveness of its individual components. In the context of high-level planning, we utilize GPT-4 [87] and Llama-2 [119], illustrating the effectiveness of our prompts across different language models. For low-level motion generation control, our baselines include MDM [117], MotionDiffuse [172], ReMoDiffuse [173], and MotionGPT [46], which span a range of text-to-motion approaches trained on HumanML3D [34] and show the generalizability of our framework. To evaluate the dynamics model, we include different design choices: (i) unconditional dynamics model which operates object dynamics independently of human motion; (ii) using human marker features as actions to the dynamics model, similar to [148]; (iii) using unprocessed human motion and object pointcloud motion as input to the dynamics model; (iv) our proposed vertex-based actions where only the contacting vertices are used for control.
4.2 Quantitative Results
In Table 1, comparing our framework to baselines with unconditional dynamics model, InterDreamer achieves better interaction quality in terms of CMD and penetration scores, showing the importance of human influence on object motion. Against methods that utilize direct raw human motion or markers for action features, our method demonstrates enhanced performance by offering more fine-grained guidance and extracting generalizable features for dynamics modeling. Tables 2 and 3 present a comparative analysis of our approach of combining high-level planning with low-level control, where we adopt various text-to-motion models against their counterparts without high-level planning on the BEHAVE and OMOMO datasets. Our approach consistently outperforms baselines. Specifically, InterDreamer exhibits superior motion quality, reflected by a lower FID, higher R-Precision, and better diversity, highlighting the benefits of incorporating our planning to reduce the distribution gap for the motion generator to generalize in the HOI synthesis task.
Figure 8: Ablation study on the dynamics model. Given the text description of “A person walks clockwise while holding a small box with left hand,” our (b) vertex-based control can synthesize consistent contacts, which (a) the baseline fails to do.
4.3 Qualitative Results
Figure 3 displays several results guided by the free-form text. Our method exhibits proficiency in interpreting the textual input and synthesizing dynamic, realistic interactions, despite the absence of training with text-interaction paired data. More importantly, as illustrated in Figure 4, we selectively use more complex sequences of interactive descriptions that are beyond the scope of the existing HOI dataset. Figure 5 further exemplifies our method that is able to generalize effectively to the CHAIRS dataset, despite our dynamics model not being trained on it. Figure 6 depicts the retrieval procedure, resulting in a diverse set of interactions that are both high-quality and semantically aligned. More experimental results and the user study are presented in Sec. D of the Appendix.